OpenStudy (blackops2_143):

What is the circumference of a circle with a radius of 0.8 m?

OpenStudy (lfghjyt):

5.03

OpenStudy (lfghjyt):

m

OpenStudy (baru):

\[circumfrance =2 \times \pi \times r\]

OpenStudy (sachintha):

Substitute the value of radius into the equation baru mentioned and simplify it.

OpenStudy (asad786):

\[2\]

OpenStudy (asad786):

2 pie r

OpenStudy (asad786):

2 times 0.8 times 3.14